Grade 12Physical Chemistry

1 gm metal M3+ was discharged by the passage of1.81 x 1023 electons.find atomic weight

akash
9 Years agoGrade 12

2 Answers

Vikas TU
9 Years ago
Moles of electrons that were being discharged are => 1.81*10^23/(6*10^23) = > 0.301 moles.
Atomic weight = > moles / weight => 0.301/1=> 0.301 g
 
Tanishq pandey
8 Years ago
For discharging of metal the reaction is :M3+ + 3e- → MThus 1 mol of the metal is deposited by 3F = 3×96500 C = 289500 C​Charge on 1.81×1023 electrons = 1.602×10-19 ×1.81​×1023 C = 2.9×104 C289500 C deposit = 1 mol of metal Thus 2.9×104 C will deposit = 1289500×2.9×104moles metal = 0.100 moles0.100 moles = 1 g metalThus atomic mass of metal = 10 g/mol